# Higher-rank GBS groups: non-positive curvature and biautomaticity

### v2

@@ 1. Introduction
In this note, we generalise the results of Leary and Minasyan to all rank $n$ generalised Baumslag–Solitar groups (or rank $n$ GBS groups for short)—fundamental groups of (finite) graphs of groups ${\mathcal{G}}$ with all vertex and edge groups isomorphic to ${\mathbb{Z}}^{n}$ for a fixed $n\in{\mathbb{N}}$. (We call such a graph of groups ${\mathcal{G}}$ a rank $n$ GBS graph of groups.) Rank 1 GBS groups are well studied, for instance from the point of view of splittings, quasi-isometries, automorphisms and commensurability [Why01, For03, For06, Lev07, Lev15, CRKZ21]. The general setting of rank $n$ GBS groups has also been studied by several authors in recent years, from the perspective of actions on hyperbolic spaces in [But22] and the perspective of separability properties in [LdGZS23].
@@
    $\pi_{1}({\mathcal{G}})$ is $\operatorname{CAT}(0)$Leftrightarrow the image of $\Delta_{\mathcal{G}}$ is conjugate in $\operatorname{GL}_{n}({\mathbb{R}})$ into $\operatorname{O}(n)$.
After writing the first version of this paper, we learnt that part (ii) of Theorem 1.1 had been independently proved by Button and was due to appear in his upcoming publication [But22].
[…]
In [LdGZS23], Lopez de Gamiz Zearra and the first author characterised when a rank $n$ GBS group is residually finite, subgroup separable and cyclic subgroup separable. In particular, they showed that $\pi_{1}({\mathcal{G}})$ is residually finite if and only if it is either an ascending HNN extension of ${\mathbb{Z}}^{n}$, or the image of $\Delta_{\mathcal{G}}$ is conjugate in $\operatorname{GL}_{n}({\mathbb{Q}})$ into $\operatorname{GL}_{n}({\mathbb{Z}})$. Consequently, if $\pi_{1}({\mathcal{G}})$ is both $\operatorname{CAT}(0)$ and residually finite, then the image of $\Delta_{\mathcal{G}}$ is conjugate into a discrete subgroup of $\operatorname{O}(n)$, so the image is finite, implying that $\pi_{1}({\mathcal{G}})$ is biautomatic. Therefore, the question of whether every residually finite $\operatorname{CAT}(0)$ group is biautomatic [LM21, Question 12.4] remains open.
@@ 2.2. Graphs of groups
Now, let ${\mathcal{G}}$ be a graph of groups, and $T\subseteq|{\mathcal{G}}|$ a maximal subtree. The fundamental group of ${\mathcal{G}}$ (with respect to $T$), denoted $\pi_{1}({\mathcal{G}},T)$, is a group with
  • generators $E^{+}{\mathcal{G}}\sqcup\bigsqcup_{v\in V{\mathcal{G}}}G_{v}$;
  • relations of the form $t_{e}\iota_{e}(g)t_{\overline{e}}=\iota_{\overline{e}}(g)$ for all $e\in E^{+}{\mathcal{G}}$ and $g\in G_{e}$ and of the form $t_{e}=1$ for all $e\in E^{+}T$.
@@ 2.3. Commensurating graphs of groups
We say a graph of groups ${\mathcal{G}}$ is commensurating if we have $[G_{i(e)}:\iota_{e}(G_{e})]<\infty$ for every $e\in E^{+}{\mathcal{G}}$—or equivalently, if the Bass–Serre tree ${\mathcal{T}}_{\mathcal{G}}$ is locally finite. In this case, it is easy to see that for any vertex $v\in V{\mathcal{G}}$ (or edge $e\in E^{+}{\mathcal{G}}$), the group $H=G_{v}$ (or $H=G_{e}$) is commensurated in $G=\pi_{1}({\mathcal{G}})$, and the map $\Phi_{G,H}$ defined above is called a modular homomorphism for ${\mathcal{G}}$ and denoted by $\Delta_{{\mathcal{G}},H}$. Moreover, every vertex and edge group $H^{\prime}$ of ${\mathcal{G}}$ shares a finite-index subgroup with $H$, implying that there exists an isomorphism $\gamma\colon\operatorname{Comm}(H)\to\operatorname{Comm}(H^{\prime})$ satisfying $\Delta_{{\mathcal{G}},H^{\prime}}=\gamma\circ\Delta_{{\mathcal{G}},H}$. Therefore, up to such an isomorphism $\gamma$, the modular homomorphism $\Delta_{{\mathcal{G}},H}$ does not depend on the choice of $H$, and we write $\Delta_{\mathcal{G}}$ for this map.
@@ 3. Biautomaticity
By [LdGZS23, Proposition 2.2], it is enough to show that $G\coloneqq\Delta_{\mathcal{G}}(\pi_{1}({\mathcal{G}}))<\operatorname{GL}_{n}({\mathbb{Q}})$ is conjugate to a subgroup of $\operatorname{GL}_{n}({\mathbb{Z}})$. But indeed, consider the lattice $H_{0}={\mathbb{Z}}^{n}<{\mathbb{Q}}^{n}$ (where by lattice we mean a subgroup isomorphic to ${\mathbb{Z}}^{n}$, i.e. a discrete subgroup of ${\mathbb{Q}}^{n}$ that spans ${\mathbb{Q}}^{n}$ over ${\mathbb{Q}}$). Then $H\coloneqq\bigcap_{A\in G}A(H_{0})$ is a lattice in ${\mathbb{Q}}^{n}$ (since any intersection of two lattices is a lattice and since $G$ is finite), and by construction, $H$ is $G$-invariant. Any linear transformation mapping a basis of $H$ (as a free abelian group) to the standard basis of ${\mathbb{Q}}^{n}$ will then conjugate $G$ to a subgroup of $\operatorname{GL}_{n}({\mathbb{Z}})$, as required. ∎
@@ 4. Non-positive curvature
[LM21, Theorem 6.4($4^{\prime}$)] Let $L$ be a finite-rank free abelian group acting properly by semisimple isometries on a CAT(0) space $X$, and let $\varphi$ be an isometry of $X$ that commensurates $L$. Then there is an inner product $\langle\cdot,\cdot\rangle_{L}$ on $L\otimes{\mathbb{Q}}$ (depending only on the action of $L$ on $X$) which is preserved by the image of $\varphi$ in $\operatorname{GL}(L\otimes{\mathbb{Q}})$ (see Subsection 2.1).

### v3

@@ 1. Introduction
In this note, we generalise the results of Leary and Minasyan to all rank $n$ generalised Baumslag–Solitar groups (or rank $n$ GBS groups for short)—fundamental groups of (finite) graphs of groups $\mathcal{G}$} with all vertex and edge groups isomorphic to ${\mathbb{Z}}^{n}$ for a fixed $n\in{\mathbb{N}}$. (We call such a graph of groups $\mathcal{G}$} a rank $n$ GBS graph of groups.) Rank 1 GBS groups are well studied, for instance from the point of view of splittings, quasi-isometries, automorphisms and commensurability [Why01, For03, For06, Lev07, Lev15, CRKZ21]. The general setting of rank $n$ GBS groups has also been studied by several authors in recent years, from the perspective of actions on hyperbolic spaces in [But25] and the perspective of separability properties in [LdGZS25].
@@
    $\pi_{1}({\mathcal{G}})$ is $\operatorname{CAT}(0)$ $\Leftrightarrow$ the image of $\Delta_{\mathcal{G}}$ is conjugate in $\operatorname{GL}_{n}({\mathbb{R}})$ into $\operatorname{O}(n)$.
After writing the first version of this paper, we learnt that part (ii) of Theorem 1.1 had been independently proved by Button and was due to appear in his upcoming publication [But25].
Note that, as a consequence of Theorem 1.1, any biautomatic rank $n$ GBS group is $\operatorname{CAT}(0)$, since any finite subgroup of $\operatorname{GL}_{n}({\mathbb{R}})$ is conjugate to a subgroup of $\operatorname{O}(n)$. Indeed, if $G<\operatorname{GL}_{n}({\mathbb{R}})$ is finite, then we can define a $G$-invariant inner product on ${\mathbb{R}}^{n}$ by setting $\langle x,y\rangle\coloneqq\sum_{A\in G}\langle Ax,Ay\rangle_{0}$, where $\langle{-},{-}\rangle_{0}$ is the standard inner product on ${\mathbb{R}}^{n}$; then any linear transformation mapping an orthonormal basis with respect to $\langle{-},{-}\rangle$ to the standard basis of ${\mathbb{R}}^{n}$ will conjugate $G$ into $\operatorname{O}(n)$.
In [LdGZS25], Lopez de Gamiz Zearra and the first author characterised when a rank $n$ GBS group is residually finite, subgroup separable and cyclic subgroup separable. In particular, they showed that $\pi_{1}({\mathcal{G}})$ is residually finite if and only if it is either an ascending HNN extension of ${\mathbb{Z}}^{n}$, or the image of $\Delta_{\mathcal{G}}$ is conjugate in $\operatorname{GL}_{n}({\mathbb{Q}})$ into $\operatorname{GL}_{n}({\mathbb{Z}})$. Consequently, if $\pi_{1}({\mathcal{G}})$ is both $\operatorname{CAT}(0)$ and residually finite, then the image of $\Delta_{\mathcal{G}}$ is conjugate into a discrete subgroup of $\operatorname{O}(n)$, so the image is finite, implying that $\pi_{1}({\mathcal{G}})$ is biautomatic. Therefore, the question of whether every residually finite $\operatorname{CAT}(0)$ group is biautomatic [LM21, Question 12.4] remains open.
@@ 2.2. Graphs of groups
Now, let $\mathcal{G}$} be a graph of groups, and $T\subseteq|{\mathcal{G}}|$ a maximal subtree. The fundamental group of $\mathcal{G}$} (with respect to $T$), denoted $\pi_{1}({\mathcal{G}},T)$, is a group obtained from the free product $_{v\in V{\mathcal{G}}}G_{v}$ by adding
  • generators $\{t_{e}\mid e\in E^{+}{\mathcal{G}}\}$;
  • relations of the form $t_{e}\iota_{e}(g)t_{\overline{e}}=\iota_{\overline{e}}(g)$ for all $e\in E^{+}{\mathcal{G}}$ and $g\in G_{e}$ (note that this includes relations $t_{e}t_{\overline{e}}=1$) and of the form $t_{e}=1$ for all $e\in E^{+}T$.
@@ 2.3. Commensurating graphs of groups
We say a graph of groups $\mathcal{G}$} is commensurating if we have $G_{i(e)}:\iota_{e}(G_{e})]<\infty$ for every $e\in E^{+}{\mathcal{G}}$—or equivalently, if the Bass–Serre tree ${\mathcal{T}}_{\mathcal{G}}$ is locally finite. In this case, it is easy to see that for any vertex $v\in V{\mathcal{G}}$ (or edge $e\in E^{+}{\mathcal{G}}$), the group $H=G_{v}$ (or $H=G_{e}$) is commensurated in $G=\pi_{1}({\mathcal{G}})$, and the map $\Phi_{G,H}$ defined above is called a modular homomorphism for $\mathcal{G}$} and denoted by $\Delta_{{\mathcal{G}},H}$. Moreover, every vertex and edge group $H^{\prime}$ of $\mathcal{G}$} shares a finite-index subgroup with $H$, implying that there exists an isomorphism $\gamma\colon\operatorname{Comm}(H)\to\operatorname{Comm}(H^{\prime})$ satisfying $\Delta_{{\mathcal{G}},H^{\prime}}=\gamma\circ\Delta_{{\mathcal{G}},H}$. Therefore, up to such an isomorphism $\gamma$, the modular homomorphism $\Delta_{{\mathcal{G}},H}$ does not depend on the choice of $H$, and we write $\Delta_{\mathcal{G}}$ for this map.
@@ 3. Biautomaticity
By [LdGZS25, Proposition 2.4], it is enough to show that $G\coloneqq\Delta_{\mathcal{G}}(\pi_{1}({\mathcal{G}}))<\operatorname{GL}_{n}({\mathbb{Q}})$ is conjugate to a subgroup of $\operatorname{GL}_{n}({\mathbb{Z}})$. But indeed, consider the lattice $H_{0}={\mathbb{Z}}^{n}<{\mathbb{Q}}^{n}$ (where by lattice we mean a subgroup isomorphic to ${\mathbb{Z}}^{n}$, i.e. a discrete subgroup of ${\mathbb{Q}}^{n}$ that spans ${\mathbb{Q}}^{n}$ over $\mathbb{Q}$}). Then $H\coloneqq\bigcap_{A\in G}A(H_{0})$ is a lattice in ${\mathbb{Q}}^{n}$ (since any intersection of two lattices is a lattice and since $G$ is finite), and by construction, $H$ is $G$-invariant. Any linear transformation mapping a basis of $H$ (as a free abelian group) to the standard basis of ${\mathbb{Q}}^{n}$ will then conjugate $G$ to a subgroup of $\operatorname{GL}_{n}({\mathbb{Z}})$, as required. ∎
@@ 4. Non-positive curvature
[LM21, Theorem 6.4($4^{\prime}$)] Let $L$ be a finite-rank free abelian group acting properly by semisimple isometries on a CAT(0) space $X$. Then there is an inner product $\langle\cdot,\cdot\rangle_{L}$ on $L\otimes{\mathbb{Q}}$, such that, for any isometry $\varphi$ of $X$ that commensurates $L$, the inner product $\langle\cdot,\cdot\rangle_{L}$ is preserved by the image of $\varphi$ in $\operatorname{GL}(L\otimes{\mathbb{Q}})$ (see Subsection 2.1).